Find the value of energy of a wave with a frequency of 8 Hz. Given: h = 6.63x1034 J f= 8Hz or 8/s E=?

1
Expert's answer
2022-01-24T08:44:58-0500

Given:

"f=8\: \rm Hz"

"h = 6.63*10^{-34}\:\rm J/s"


The energy of photon

"E=hf=\rm 6.63*10^{-34}\:\rm J/s*8\: Hz=5.3*10^{-33}\: J"